OpenStudy (anonymous):

The product of 4 and a number is 18 more than twice the number. What is the number?

OpenStudy (anonymous):

i think 9

OpenStudy (cwrw238):

4x = 2x + 18 2x = 18

OpenStudy (rhystic):

so product means multiply which gives 4*x = 18 + 2x -> 2x = 18 -> x =9

OpenStudy (cwrw238):

so x = ?

OpenStudy (anonymous):

so 9 is

OpenStudy (anonymous):

right

OpenStudy (anonymous):

4 x - 18 = 2 x // Solve x = 9

OpenStudy (cwrw238):

yes 9 is right